i need to solve for y and x . i dont quit understand it.here is the problem 7y+15=3x ,15=3x+2y . this is one problem the teacher said i have to combine them.to find y and x
Found 2 solutions by stanbon, Leaf W.:
Answer by stanbon(75887)

7y+15=3x
15=3x+2y
---
Rearrange:
3x = 7y+15
3x = 15-2y
------------
Substitute and solve for "y":
7y+15 = 15-2y
9y = 0
y = 0
-----
solve for "x":
3x = 7y+15
3x = 7*0+15
3x = 15
x = 5
---------
x = 5 and y = 0
======================
cheers,
Stan H.
================

Answer by Leaf W.(135)

1. Put them in the same format. Let us decide on the format that the first one, 7y + 15 = 3x, is in. To put the second one in this format, subtract 2y from both sides of the equation:
-2y + 15 = 3x.
2. Make it so that one of the variables (x or y) in the equations would cancel out if the two equations were to be added or subtracted from one another. Since 3x and 3x would cancel out if subtracted, you can skip this step.
3. Subtract the two equations so that the 3x is canceled out.
7y + 15 = 3x
-(-2y + 15 = 3x)
This would be the same as
7y + 15 = 3x
+ 2y - 15 = -3x
--> 9y = 0
4. Solve for the value of the remaining (y) variable.
9y = 0
Divide both sides by 9.
y = 0
5. Plug this value in for y in one of the equations (it does not matter which one).
7y + 15 = 3x
7(0) + 15 = 3x
0 + 15 = 3x
15 = 3x
5 = x
6. For your answer, plug in your values for x and y into coordinate form (x,y). ==> (5,0)
Good luck! =)